OpenStudy (anonymous):

WHAT VALUE OF K WOULD MAKE THE GRAPH OF 6X+2Y=K AND THE GRAPH OF 4X+KY=3 PARALLEL

OpenStudy (anonymous):

4/3?

OpenStudy (anonymous):

Check it.

OpenStudy (anonymous):

How do I Check it ? do I plug something in?

OpenStudy (anonymous):

OK, let me check it, then I will explain.

OpenStudy (anonymous):

thank you and just let me know.

OpenStudy (anonymous):

Ok, it seems right.

OpenStudy (radar):

I think you nailed it estudier

OpenStudy (anonymous):

Do you know how to get the slope of a line?

OpenStudy (anonymous):

no I don't I was just transferred into a new school and didnt have much time to catch up! that's why I need to know like how did you get 4/3? so if i see it again I'll know what to do.

OpenStudy (anonymous):

OK, if u have a line in the form ax + by +c = 0 (c can be on the other side of =) Then the gradient or slope is -a/b, follow me?

OpenStudy (anonymous):

OKAY IM WITH YOU SO FAR !

OpenStudy (anonymous):

For the lines to be parallel, the slopes must be the same. So put slope 1 = slope 2 and solve for K.

OpenStudy (anonymous):

SO WOULD IT LOOK LIKE THS 6X+2Y=5+4X+KY+3? STILL A LITTLE CONFUSED

OpenStudy (anonymous):

I MEANT =3

OpenStudy (anonymous):

The slope is -a/b, u said u understood. What is the slope of the first line?

OpenStudy (anonymous):

THE EQUATION IS 6X+2Y=5 SO WOULD THE SLOPE BE 6

OpenStudy (anonymous):

ax + by = c 6x + 2y = 5 Slope is -a/b, compare the two and tell me the slope.

OpenStudy (anonymous):

So the a is 6....yes?

OpenStudy (anonymous):

YEAH!

OpenStudy (anonymous):

and the b?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right. So -a/b is?

OpenStudy (anonymous):

-6/2

OpenStudy (anonymous):

IS THAT RIGHT

OpenStudy (anonymous):

Ok. -3. Now do the same for the other line.

OpenStudy (anonymous):

OKAY HOLD ON !

OpenStudy (anonymous):

THE SECOND ONE LOOK LIKE THIS 4X+KY=3 SO 4 WOULD BE A AND SINCE IT ISNT A VALUE FOR KY WOULD IT BE 0 OR 1

OpenStudy (anonymous):

It is a value, the value is K.

OpenStudy (anonymous):

So the slope is..?

OpenStudy (anonymous):

SO THE SLOPE IS K....RIGHT?

OpenStudy (anonymous):

Do it exactly the same way we did the first one.

OpenStudy (anonymous):

a is...?

OpenStudy (anonymous):

4 AND B IS K SO WOULD IT BE -4/3 K

OpenStudy (anonymous):

No....-4/K

OpenStudy (anonymous):

OKAY THANKS ! I APPRECIATE IT !

OpenStudy (anonymous):

Not finished yet.

OpenStudy (anonymous):

OKAY WELL LETS FINISH!

OpenStudy (anonymous):

You have calculated slope 1 as -3 and slope 2 as -4/K. For the lines to be // the 2 slopes must be equal so -3 = -4/K. So what's K?

OpenStudy (anonymous):

SO WHAT DO I PUT IN THE CALCULATOR?

OpenStudy (anonymous):

U don't need a calculator to do that!

OpenStudy (anonymous):

SO DO I DIVIDE ON BOTH SIDES?

OpenStudy (anonymous):

Multiply both sides by K.

OpenStudy (anonymous):

OKAY HOLD ON

OpenStudy (anonymous):

I think u have a lot of work to do to catch up, this last part is very simple math.

OpenStudy (anonymous):

-3 = -4/K Multiply by K -3K = -4 Divide by -3 K = -4/-3 = 4/3.

OpenStudy (anonymous):

Which is the answer to the question.

OpenStudy (radar):

Good job estudier

OpenStudy (anonymous):

Thanks!:-)

OpenStudy (anonymous):

Good luck with your studies Bhen--

OpenStudy (anonymous):

THANKS SO MUCH I JUST HAD TO WRITE DOWN EVERYTHING STEP BY STEP SO I CAN KNOW NEXT TIME ..YOU ARE VERY HELPFUL AND I HOPE TO GET HELP FROM U IN THE NEAR FUTURE!!